OpenStudy (anonymous):

how to solve the Differential equation: y'+2xy=2x , y(0)=2 how to solve the Differential equation: y'+2xy=2x , y(0)=2 @Mathematics

OpenStudy (anonymous):

do you know Laplace Transformation?

OpenStudy (anonymous):

turn y' into dy/dx first

OpenStudy (anonymous):

no, im not sure how to do the laplace transformation.

OpenStudy (anonymous):

ok.. dy/dx-2xy=2x (1) The differential equation without the second member is the following. dy/dx-2xy=0 dy/dx=2xy dy/y=2xdx ln|y|=x²+lnC y=Ce^x² Let's note dy/dx=y'. We can use the method of the variation of the constant. y'=C'e^x²+2xCe^x² (1) is written as follows. C'e^x²+2xCe^x²-2xCe^x²=2x C'e^x²=2x C'=2xe^-x² C=-e^-x²+K y=Ce^x² y=-1+Ke^x² The solution of the differential equation is y=Ke^x²-1

OpenStudy (stacey):

Since this is in the form y' + p(x)y=g(x), it is a linear differential equation and can be solved using the integrating factor \[u(x)=e ^{\int\limits_{}^{}p(x)dx}\]

OpenStudy (stacey):

Since p(x)=2x, the integrating factor is \[u(x)=e ^{x ^{2}}\]

OpenStudy (stacey):

Multiplying by the integrating factor, our equation changes to\[e ^{x ^{2}}y' + 2xe ^{x ^{2}}y = 2xe ^{x ^{2}}\]

OpenStudy (stacey):

We can take the integral of both sides. Notice that the left side looks like what occurs when we take the derivative using the product rule, so when we integrate that side we end up with a product. \[e ^{x ^{2}}y=e^{x ^{2}}+C\]

OpenStudy (stacey):

Divide to get y by itself and we have\[y=1+Ce^{-x ^{2}}\]

OpenStudy (stacey):

Using the initial conditions of y=2 when x=0, we can solve for C and get C=1.

OpenStudy (anonymous):

thank you! :)

OpenStudy (stacey):

You're welcome.

OpenStudy (anonymous):

y'+2xy=2x , y(0)=2 dy/dx+2xy-2x =0 dy+2xydx-2xdx =0 dy+2xdx(y-1) =0 dy=-2xdx(y-1) dy/(y-1) =-2xdx integrating both sides gives

OpenStudy (anonymous):

ln(y-1)=-x^2 +c whenx=0,y=2 c=0 y-1=e^(-x^2) y=1+e^(-x^2) ans

OpenStudy (anonymous):

this is the solution to the DE y'+2xy=2x , y(0)=2

OpenStudy (stacey):

You did it by separating them. I didn't see that. :-)